A patient weights 146 lbs and is to recieve a dosage of 45.0 mg per kg of body weight. the drug is supplied as solution that contains 25.0 mg of drug per mL of solution. what volume of drug should the pateint recieve?


1
Expert's answer
2021-09-21T01:57:04-0400

146/2.205 = 66.224 kilograms

45× 66.224 = 2980.1mg

2980.1/25 = 119.2032ml